# Massimiliano Ciranni1 These authors contributed equally to this work.    Vito Paolo Pastore1,2∗    Roberto Di Via1∗    Enzo Tartaglione3    Francesca Odone1    Vittorio Murino2,4

### v2

@@ 4.3.2 Debiasing accuracy
Table 2: Results of DDB Recipes I and II on realistic datasets. Ticks () indicate that the method is bias-unsupervised. We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
Regarding the considered real-world datasets, we can highlight a similar trend in performance, with DDB’s approaches outperforming the current state of the art (see Table 2). Notably, in Waterbirds, WGA is beating the G-DRO supervised (91.56% vs. 91.40% of G-DRO [8]) and all other unsupervised methods’ accuracies. DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]). In BFFHQ, DDB-I and DDB-II obtain accuracies of 74.67% and 70.93%, respectively, as compared with 69.3%, the best state-of-the-art results to date (MoDAD [3]), improving of 6.34 and 2.6 percentage points, respectively.
@@ 4.4 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Suppl. Material. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data. Biased data generation break point. We aim here to empirically estimate the degree of correlation $\rho$ in the training set leading the CDPM to predominantly learn bias-aligned samples. Starting from the original real images, we build ad hoc controlled training sets such that the number of total samples is kept fixed ($4060$ images, balanced among the classes), while the ratio between aligned and conflicting samples varies. Specifically, we increase the proportion of bias-conflicting from $0.05$ to $0.30$ ($\rho=0.950-0.700$), training a CDPM with CFG on each examined ratio. Then, we generate $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples. This aims to assess to which extent the bias ratio $\rho$ permits a correct approximation of the biased data distribution, allowing a proper training of the auxiliary model, while investigating the impact on debiasing performance with our Recipe I. The obtained results are highlighted in Table 3. As expected, for $\rho=0.950$, a qualitative analysis reveals very few conflicting generated images. This is still substantially true up to $\rho=0.800$, where we find $8.80\%$ bias-conflicting samples in the generated images, reflected by good debiasing performance ($88.69\%$ and $86.52\%$ for $\rho=0.900$ and $\rho=0.800$). When $\rho=0.700$, as expected, despite the CDPM still amplifying bias-aligned samples, a relevant percentage of bias-conflicting generations emerges ($17.3\%$), impacting final performance.
Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
WGA $\bf 90.81$ $88.69$ $86.52$ $67.22$
Table 3: Ablation study on the impact of $\rho$ on correlation bias in synthetic images and WGA performance in Waterbirds.
Bias amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class. Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier. From Table 4 (left), we observe that even 100 synthetic images are enough for good debiasing performance, despite not being competitive with top accuracy scores. As we increase the number of training images, WGA improves, reaching the best performance for $1000$ images per class
@@
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Tables 1 and 2, DDB-I (w/ plain err. set, last row in the tables) report our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios ($\rho=0.980$ and $\rho=0.995$).
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present. This ablation study aims to assess DDB from this perspective. To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on the original CIFAR-10 dataset [44], comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
@@ 5 Conclusions
Limitations
DDB’s main limitations rely on the well-known high computational complexity of diffusion models. As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM. Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM. Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
@@
Ablation Studies with DDB’s Recipe II
@@
Biased data generation break point.
Using the same protocol employed for Recipe I, we want to study the impact of the input training set for the CDPM on the whole pipeline. In the main paper, we describe the process of building ad-hoc training sets (where all have a total of $4060$ images, balanced among the classes, but a different ratio between aligned and conflicting samples) from Waterbirds, and the subsequent generation of $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples. Using the same pools of generated images, we train a dedicated Bias Amplifier and use it in our Recipe II. The obtained results, outlined in Table 5, resemble what is shown in the main paper for Recipe I (Table 3). As expected, when the input training set for the CDPM has $\rho=0.900$, resulting generations are still substantially dominated by aligned samples, allowing for training an effective Bias Amplifier. This is reflected by the final WGA ($90.03\%$). A similar behavior can be found for $\rho=0.800$ (WGA = $89.53\%$). As it was for Recipe I, $\rho=0.700$ results in a sensibly less effective Bias Amplifier due to the relevant percentage of conflicting samples among generated images. As a consequence, the final WGA is negatively affected ($73.52\%$).
Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
WGA $\bf 91.56$ $90.03$ $89.53$ $73.52\%$
Table 5: Ablation on the number of training images for the Bias Amplifier, in terms of resulting WGA on Waterbirds, when using DDB’s Recipe II.
@@
Similarly to what is shown in Sec. 4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance. Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on the original test set of CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
Full Implementation Details of Recipe I (two-step)
@@
For the hyperparameters of Recipe I, we use for each dataset the following configurations. Corrupted CIFAR-10: regardless of the specific $\rho$, we set batch size = 32, learning rate = $1\times e^{-4}$, weight decay = 0.01, and training is set to a total of 150 epochs. BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs. The GDRO optimization is always run with the --robust and --reweight_groups flags.
Full Implementation Details of Recipe II (end-to-end)
@@
For Corrupted CIFAR-10 and BAR, we use the same parameters reported in [6]. Corrupted CIFAR-10: Regardless of the specific $\rho$, we set batch size = 256, learning rate = $5\times e^{-4}$, weight decay = $0.01$, and training is set to a total of 50 epochs. BAR: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For the real-world datasets, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
Additional Biased Generations
@@
In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, comprising three real-world and one synthetic dataset, as produced by our CDPM. With CIFAR10-C (Figure 6), the model proficiently synthesizes classes with specific corruptions, evidencing its ability to replicate synthetic bias patterns. In the Waterbirds dataset (Figure 7), the model adeptly captures and replicates the correlation between bird species and their background. In the BFFHQ dataset (Figure 8), it effectively reflects demographic biases by mirroring gender-appearance correlations. The results from the BAR dataset (Figure 9) exhibit the model’s adeptness in handling multiple bias patterns, preserving typical environmental contexts for different actions. Overall, the model generates high-quality samples that are fidelity-biased across all datasets. Our research confirms that the model exhibits bias-learning behavior, as it not only learns but also amplifies existing biases within diverse datasets. We leverage this characteristic to produce synthetic biased data aimed at enhancing the robustness of debiasing techniques.
Figure 6: Synthetic bias-aligned image generations for each class of the CIFAR10-C dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
@@
Figure 7: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Figure 8: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
@@
Figure 9: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Bias Validation via Captions in Synthetic Images

### v3

@@ 4.3.2 Classification accuracy
Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () indicate that the method is bias-unsupervised. We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1). Notably, in Waterbirds, DDB’s WGA is higher than G-DRO supervised (91.56% vs. 91.40% of G-DRO [8]) and all the other unsupervised methods’ ones. DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to ETF-Debias [43] (73.60%), the best state-of-the-art results to date, and provides a boost of 7.45% over the recent DeNetDM [45].
@@ 4.4 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Suppl. Material. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
Bias amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class. Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier. From Table 2 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores. As we increase the number of training images, WGA improves, reaching the best performance for $2000$ images per class.
@@
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Table 1, the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$) (e.g., Waterbirds), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like BFFHQ ($\rho=0.995$).
Bias-Identification Accuracy To frame the goodness of our Bias Amplifier in correctly identifying aligned and conflicting training samples, we provide the accuracy of such process when considering samples misclassified by the BA. Specifically, we report a bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ. Our very high accuracy supports the effectiveness of using the proposed protocol, which is not dependent from any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present. This ablation study aims to assess DDB from this perspective. To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  [46], comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
@@ 5 Conclusions
Limitations. DDB’s main limitations rely on the well-known high computational complexity of diffusion models. As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM. Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM. Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
@@
Appendix A Ablation Studies with DDB’s Recipe II
@@
Similarly to what is shown in Sec. 4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance. Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
Appendix B Full Implementation Details of Recipe I (two-step)
@@
For the hyperparameters of Recipe I, we use for each dataset the following configurations. BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs. The GDRO optimization is always run with the --robust and --reweight_groups flags.
Appendix C Full Implementation Details of Recipe II (end-to-end)
@@
For BAR, we use the same parameters reported in [6]: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
Appendix D Additional Biased Generations
@@
In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, as produced by our CDPM. In the Waterbirds dataset (Figure 6), the model adeptly captures and replicates the correlation between bird species and their background. In the BFFHQ dataset (Figure 9), it effectively reflects demographic biases by mirroring gender-appearance correlations. The results from the BAR dataset (Figure 8) exhibit the model’s adeptness in handling multiple bias patterns, preserving typical environmental contexts for different actions. Finally, the ImageNet-9 dataset (Figure 7) stereotypes the various classes and their backgrounds well. Overall, the model generates high-quality samples that are fidelity-biased across all datasets. Our research confirms that the model exhibits bias-learning behavior, as it not only learns but also amplifies existing biases within diverse datasets. We leverage this characteristic to produce synthetic biased data aimed at enhancing the robustness of debiasing techniques.
@@
Figure 6: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
@@
Figure 7: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Figure 8: Synthetic bias-aligned image generations for each class of the ImageNet-9 dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Figure 9: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Appendix E Bias Validation via Captions in Synthetic Images